Consider the hypotheses shown below. Given that

x overbar =50​, σ=11​, n=36​, α=0.01, complete parts a and b.

H0: mu greater than or equal to  47

H1: mu less than or equal to 47

a. What conclusion should be​ drawn?

b. Determine the​ p-value for this test.
Answer #1

To Test :-
H0 :- µ > 47
H1 :- µ <= 47

Part a)

Test Statistic :-
Z = ( X - µ ) / ( σ / √(n))
Z = ( 50 - 47 ) / ( 11 / √( 36 ))
Z = 1.64


Test Criteria :-
Reject null hypothesis if Z < -Z(α)
Critical value Z(α) = Z(0.01) = 2.326
Z > -Z(α) = 1.64 > -2.326
Result :- Fail to reject null hypothesis

Part b)

P value = P ( Z < 1.6364 ) = 0.9491

Looking for the value Z = 1.64 in standard normal table to find the P value.